OpenStudy (anonymous):

What is the vertex of the function f(x)=2x^2+4x-16. I think you start with (-b/2a, f (-b/2a))

hero (hero):

@raindance, seems like a good start to me.

hero (hero):

Have you attempted to figure out the vertex yet?

OpenStudy (anonymous):

Yes. I got -1,-18

hero (hero):

Would you mind showing your work please?

OpenStudy (anonymous):

-4-2*2 = -1 2*-1^2-4-16=-18

hero (hero):

\[-\frac{b}{2a} = -\frac{4}{2(2)} = -1\] \[f(-\frac{b}{2a}) = f(-1) = 2(-1)^2 + 4(-1) - 16 = 2 - 4 - 16 = -18\] Correct

OpenStudy (anonymous):

I have one more question. Can you explain how to find the x-intercepts of f(x)=2x^2-10x+12.

OpenStudy (anonymous):

@Hero

hero (hero):

The x-intercepts of the graph of f(x) = 2x^2 - 10x + 12 occur where f(x) = 0. So... Set f(x) = 0, then solve for x: 0 = 2x^2 - 10x + 12

hero (hero):

In other words, solve the quadratic equation for x.

OpenStudy (anonymous):

-12=2X^2-10X I'm not sure hwat to do with the square...

hero (hero):

Actually, you should begin by factoring out the 2 on the right side: 0 = 2(x^2 - 5x + 6) Then divide both sides by 2: 0 = x^2 - 5x + 6 Now you have to factor x^2 - 5x + 6

OpenStudy (anonymous):

I'm having a bit of a bran freeze is that finding two numbers that add to five and multiply to six?

hero (hero):

That's part of the factoring process, yes :) Find two numbers that add to get -5 and multiply to get 6.

OpenStudy (anonymous):

-2 and -3

hero (hero):

For the next steps, You could replace 5 with 2 + 3 as follows: 0 = x^2 - (2 + 3)x + 6 Then distribute -x as follows: 0 = x^2 -2x -3x + 6 Then factor the first two terms and the last two terms on the right side: 0 = x(x - 2) -3(x - 2) And finally, factor out the remaining common term (x - 2): 0 = (x - 2)(x - 3)

hero (hero):

Afterwards, you would use the zero product property to continue solving for x.

OpenStudy (anonymous):

x-2=0 x=2 x-3=0 x=3